# The Challenge of Using LLMs to Simulate Human Behavior: A Causal Inference Perspective

### v2

@@ 3 Potential Solution: Controlling for Covariates
When working with real-world data, a common approach for causal identification is to control for covariates. If researchers have controlled for all confounding covariates that affect both the treatment and outcome, the conditional independence assumption, or ignorability, is satisfied. If sufficient variation remains after controlling for covariates, then researchers can identify the causal effect of interest. In LLM-simulated experiments, a natural way to control for covariates is adding detailed information about individuals and their environment in the prompt. These covariates can range from demographic information, location, preferences, budget, etc. We discuss other approaches including RAG and fine-tuning in the Section 4.3.
Argyle et al., (2023) introduced an approach for controlling for such covariates. They advocate for running studies on “silicon samples” of digital personas, which in their case were described by demographic variables and political beliefs that were drawn to align with nationally representative samples. This approach has been picked up by the industry and commercialized by several firms (e.g., Synthetic Users, Evidenza). Figure 3 shows that on average, controlling for covariates, such as demographic variables does indeed improve the results in our case compared to not specifying any detail. However, it does not fully resolve the issue. Appendix D describes the detailed prompts and covariates used, which involve the creating of 500 personas for each {product,category} pair.
@@
An important consideration is the stage at which covariates are controlled for. When working with human data, empirical researchers often control for product characteristics and rich user history to address endogeneity during data analysis, after the data is already collected. Hence analyzing data does not change the original data generating process that already took place. In contrast, adding covariates in the prompt in LLM simulation occurs at the data generation stage. Consequently, the additional covariates can change the factors considered by a simulated individual when making decisions.
@@
For example, in the context of demand estimation, one can include contextual information in the prompt such as the price of competing products in the category, which can be used as a proxy for the type of stores that the customer is visiting and the competitive environment that the focal product is facing. However, while adding contextual variables such as competing prices in the LLM simulation may prevent the LLM from using the price treatment to infer these contextual variables, adding them may also lead the LLM-simulated customer to put a heavier emphasis on these factors when making decisions. Figure 4 shows that the demand curve (elicited by simulating the choices of 500 personas) is no longer smooth after fixing such contextual variables, [^5] We fix the competing price to be the same as the regular price of the focal product. and the simulated customers’ purchase decisions follow a step function: customers will purchase if and only if the price is lower than or equal to the competing price. Appendix E describes the detailed prompts. [^6] One possible way to fix the non-smoothness of the demand curve is to simulate a distribution of contextual variables, each with difference past and competing prices. However, this is no longer useful for a particular context, e.g., a particular store, because the simulated context deviates from the targeted context. In a real-world scenario, we should expect that even in the absence of heterogeneity in context (e.g., in a given store in a given day), there should still exists heterogeneity in consumers, such that the demand curve would be smoother.
@@
Figure 4: Purchase probability of Coca-Cola when controlling for demographic variables and competing price
@@
Such phenomenon in which decision makers place more emphasis on certain aspects of a choice set or scenario than they would in real-world decisions is known in the consumer behavior literature as focalism (Schkade and Kahneman,, 1998; Wilson et al.,, 2000), and constitutes a threat to ecological validity. For example, in the context of conjoint analysis, Bedi and Reibstein, (2021) argue that when minor attributes are presented together with major ones, they tend to receive undue attention from respondents (due to focalism), leading to upward biased estimates of the importance of these minor features. While it has been documented that LLM is prone to focalism (Cheng et al.,, 2023), it is important to emphasize that focalism is not an issue that is unique to LLM experiments.
While traditional experiments with humans also run the risk of failing to simulate realistic scenarios, the nature of this risk differs. When running experiments with humans, using a blind design with realistic scenarios typically ensures ecological validity without introducing concerns of confoundedness. On the other hand, with LLM simulations, a blind design raises confoundedness concerns. Adressing confoundedness by adding covariates requires adding covariates in the data generation stage of the study, which raises yet another concern, focalism, which in turn reduces ecological validity. In other words, LLM simulations pose a unique dilemma to researchers. Improving ecological validity may require leaving unspecified details in the prompt, which runs the risk of introducing confoundedness. On the other hand, including extensive detail may reduce confounding, but at the expense of ecological validity due to focalism.
4 Theoretical Framework
So far we have demonstrated one empirical example with one specific version of an LLM. A natural question is whether the issue we have uncovered applies generally to all LLMs. This section presents a theoretical framework demonstrating that the challenges identified in this paper stems from ambiguous prompts rather than fundamental limitations of the models themselves. Therefore, the solution to this specific challenge lies not only in developing more advanced LLMs, but also in developing clearer prompting strategies.
4.1 Blinding Generates Ambiguous Prompts
To demonstrate how ambiguity arises in LLM prompts, we begin by examining a simple prompt $p$ under a blinded simulation design that illustrates the fundamental problem:
Prompt 3: A simple ambiguous prompt Please consider the following product category: {category}. Suppose you are in a grocery store, and you see the following product in that category: {product}. The product is priced at: {price}. Would you or would you not purchase the product?
The design being blind, incomplete information is provided about the distribution from which the various variables in the prompt are drawn. As a result, this prompt is ambiguous. It allows for at least two interpretations:
The answers to these two questions are likely to be different. Using the do-operator (Pearl,, 2009) to distinguish mere observations and active intervention, and using $Y$ to represent $Purchase$ and $D$ to represent Price, the answer to $q_{1}$ can be represented as $a_{1}:P(Y|do(D))$ while the answer to $q_{2}$ can be represented as $a_{2}:P(Y|D)$.
$Y=\beta_{M}D+U$
where $\beta_{M}$ represents the causal effect of $D$ on $Y$ through a set of realistic mechanism $M$, and $U$ represents unobserved factors. In a well-designed experiment, we have $Cov(U,D)=0$, making it possible to identify $\beta_{M}$ and estimate $P(Y|do(D))$ from data on $(Y,D)$ alone. But under a blinded design with an ambiguous prompt, the LLM does not know it should mimic an experimental environment. That is, the LLM does not have enough information to draw $U$ from the appropriate distribution (literally or figuratively - we do not attempt to model how the LLM makes inference but rather present a conceptual model). If the model is mostly trained on observational data, it is mostly likely going to mimic observational data patterns, where $Cov(U,D)\neq 0$. As a result, the LLM’s responses will reflect confounded relationships rather than the intended causal effects. Figure 5 summarizes two possible DGPs mimicked by the LLM under this simple ambiguous prompt.
$D$$M$$Y$$U$(a) DGP for Experimental $D$$M$$Y$$U$(b) DGP for Observational Data
Figure 5: Possible DGPs mimicked by LLM simulations under ambiguous simple prompts
As discussed in Section 3, one might attempt to address this issue by adding more variables $X$ to the prompt. In principle, if $Cov(U,D|X)=0$, then controlling for $X$ can help restore identification. However, simply specifying covariates in the LLM prompt does more than condition on them statistically. Including $X$ may cause the simulation to alter the underlying decision-making mechanisms $M$ and shift toward a new set of mechanisms $M^{\prime}$ and produce an environment that differs substantially from the original scenario – the estimated parameter $\beta_{M^{\prime}}$ may have limited ecological validity, since it no longer reflects the realistic mechanisms $M$ that were originally of interest (as we illustrated in Figure 4). Figure 6 summarizes two possible DGPs mimicked by the LLM under this more complicated ambiguous prompt.
[…]
Figure 6: Potential DGPs mimicked by LLM simulations when controlling for additional variables
These graphs reveal a deeper challenge: adding variables to the prompt can actually increase rather than decrease ambiguity from a causal perspective. To illustrate this multiplication of possible interpretations in practice, consider the following detailed example:
This prompt, while detailed, is even more ambiguous. It permits more valid causal interpretations:
  1. 1.
[…]
  2. 2.
[…]
  3. 3.
[…]
4.2 Impossibility Theorem for Ambiguous Prompting Strategies
Having demonstrated how ambiguity can arise in practice, we now formalize these insights into a theoretical framework that explains why such ambiguity poses fundamental challenges for LLM evaluation and development.
Definition 4.1 (Functional Representation of a Large Language Model).
A large language model may be represented by:
$f:\mathcal{P}\rightarrow\Delta(\mathcal{Y}),$
where $\mathcal{P}$ is the set of all possible prompts, $\mathcal{Y}$ is the set of all possible outputs or responses, and $\Delta(\mathcal{Y})$ represents the set of all probability distributions over $\mathcal{Y}$. For any given prompt $p\in\mathcal{P}$, the model $f$ outputs a probability distribution $f(p)$ over potential responses $y\in\mathcal{Y}$.
[…]
@@ 4.2 Impossibility Theorem for Ambiguous Prompting Strategies
While prompting strategies can be refined empirically using validation datasets, developing a theoretical framework is essential for understanding when and why these strategies will generalize beyond the validation set. When ground truth data is available, researchers can experiment with different prompts and even fine-tune LLMs to improve performance on specific questions. However, without theoretical guidance, such empirical refinements may not yield strategies that reliably generalize to novel scenarios. This raises a fundamental question: what characteristics make a prompting strategy robust and reliable for out-of-distribution inference? We identify a particular class of prompting strategies to avoid:
@@
Such ambiguous prompting strategies should be avoided because no LLM models, no matter how advanced they are, can answer all questions correctly given such prompting strategy:
Theorem 4.1.
Proof.
4.3 Implications for model development using fine-tuning or RAG
Theorem 4.1 reveals how ambiguous prompting strategies create limitations when developing and deploying LLMs, even with techniques like fine-tuning or retrieval-augmented generation (RAG). Suppose developers leverage additional training data $(q_{1},a_{1})$ with a prompting strategy where $p=s(q_{1})$. If this prompt is ambiguous such that $p=s(q_{2})$ for some other question $q_{2}$ with correct answer $a_{2}\neq a_{1}$, then any fine-tuning that optimizes the model’s accuracy on $q_{1}$ necessarily sacrifices its accuracy on $q_{2}$. This trade-off is especially problematic given uncertainty about future applications - users may interpret the same prompts in different ways which developers cannot anticipate during model development.
Consider a concrete example where developers have access to an experimental dataset that randomizes Coca-Cola’s price while keeping Pepsi’s price at business-as-usual levels. If developers fine-tune an LLM using a blinded design that simply describes purchase decisions at different Coca-Cola prices without revealing their experimental nature, they create an ambiguous prompt that permits both experimental and observational interpretations. While the fine-tuned model may excel at predicting purchases in scenarios matching the training data ($q_{1}$), this capability is less valuable since the experimental data already directly answers such questions. The more important question is whether the model can learn the causal relationship between price and purchases, and can generalize to answer novel questions ($q_{2}$) - like what happens when Pepsi’s price changes while Coca-Cola’s price remains at business-as-usual levels, or when both prices increase simultaneously. Such generalization proves challenging theoretically.
First, under ambiguous prompts, fine-tuning based on experimental data provides no guarantee that the model will learn causal relationships. Suppose that the data generating process for the experimental dataset follows $Y=\beta_{M}D+U$, where $Y$ represents the purchase decision, $D$ represents the price of Coca-Cola, and $U$ represents everything else. The fine-tuned model may implicitly learn a regression coefficient $\widehat{b}=\frac{Cov(Y,D)}{Var(D)}=\beta_{M}+\frac{Cov(U,D)}{Var(D)}$. However, since the fine-tuned model builds on the base model, it can maintain a prior that $\widehat{\frac{Cov(U,D)}{Var(D)}}=\rho\neq 0$ and hence will not interpret the regression coefficient $\widehat{b}$ as causal. After accounting for this prior, the model may even incorrectly assume that the causal effect is $\widehat{\beta}_{M}=\widehat{b}-\rho$.
Furthermore, even if the model somehow updates its assumptions correctly to handle one specific type of question ($q_{1}$), it may fail to generalize to other closely related but distinct questions ($q_{2}$) where different assumptions are needed. For example, suppose the model learns from experimental data where Coca-Cola’s price was randomized while Pepsi’s price remained business-as-usual. Through fine-tuning, the model might correctly learn to treat Coca-Cola’s price as independent of other unobserved market factors when analyzing that specific scenario ($q_{1}$). However, when asked to predict what happens when Pepsi’s price is set to another level while Coca-Cola’s price remain business-as-usual ($q_{2}$), the model might incorrectly maintain this assumption of independence, when in fact Coca-Cola’s business-as-usual price would likely be correlated with other market factors in this new scenario. This limitation is avoidable if the experimental nature of the data is explicitly communicated to the model during both training and evaluation through less ambiguous prompts.
While the above example focuses on a case with one experimental dataset, the challenge persists even when using diverse experimental datasets for fine-tuning or retrieval-augmented generation (RAG). With RAG, this challenge manifests at both the development and deployment stages. During model development, the reference datasets should be annotated with metadata specifying which variables were randomized treatments versus correlated covariates. During retrieval, prompts should unambiguously specify the causal question to ensure the model draws from experiments with matching causal structures rather than just similar surface features. Without such explicit specification of the causal framework, neither fine-tuning nor RAG can reliably learn to distinguish between different types of variation in the data.
To summarize, our theoretical framework proves that no general-purpose model can accurately simulate all counterfactual scenarios under an ambiguous prompting strategy. This statement should not be interpreted as suggesting that fine-tuning or RAG are not valuable tools - these techniques can substantially enhance model performance by incorporating high-quality training data and reducing various forms of bias. Rather, it demonstrates that while better training data and advanced techniques are important for improving simulation accuracy, they are complementary to, rather than substitutes for, clear and unambiguous prompt design. While obtaining better data for validation, fine-tuning, or bias-correction can be resource-intensive, implementing unambiguous simulation designs may be achieved at no additional cost. The optimal approach combines both strategies: leveraging advanced training techniques while ensuring an unambiguous prompting strategy.
5 Unblinding as a Potential Solution
Our theoretical framework points toward a promising direction: designing unambiguous prompts by explicitly communicating the experimental design in the prompt, which by definition entails unblinding. Here we discuss several possible simulation designs that leverage unblinding.
One approach maintains the between-prompt structure where each prompt simulates one experimental condition, but explicitly reveals the full experimental design in each prompt. This allows generating multiple responses to form a distribution for each condition while maintaining the independence of observations across prompts. Another possibility is to use a within-prompt design where multiple experimental conditions are presented together in a single prompt with full information on the experimental design. This approach is more efficient computationally since it obtains responses for multiple conditions in one call to the model. However, showing multiple conditions together may induce artificial consistency in the responses. [^8] It is worth noting that within-prompt designs are not necessarily unblinded - a prompt could present multiple conditions (e.g., different prices) without explaining how those conditions were selected or distributed. Therefore, merely switching to within-prompt experiments is not sufficient to resolve prompt ambiguity. An even simpler approach is an unblind within-prompt aggregate design that directly asks the LLM to predict the aggregate results of the full experiment. This represents the greatest departure from traditional experimental protocols with human subjects.
Comparing the relative performance of these different approaches is beyond the scope of this paper. Unlike our analysis of blinding - where we demonstrated fundamental limitations that will persist across model generations - the effectiveness of different unblinded designs may vary substantially across models and contexts. The relative performance one may find now may no longer hold in the future, especially if future models are trained or fine-tuned based on these prompts. Due to this reason, this section only focuses on demonstrating the feasibility of the unblinded design through the simplest approach of directly eliciting aggregate predictions of the results of the experiment.
For this demonstration, we directly ask the LLM to predict aggregate experimental results across randomly assigned prices. Figure 7 shows that while this strategy’s predictions still deviate from actual consumer behavior, it does outperform the other potential solutions tested in this paper. Appendix F documents the detailed prompts.
Figure 7: Unblinded design asking directly for aggregate predictions
This approach is by no means perfect. In addition to the fact that the result still deviates from the ground truth, Figure F.1 shows that the results are also sensitive to the experimental range provided in the prompt: when the demand curve is built by combining two experiments that vary prices from [0%, 100%] and from [100%, 200%] of the regular price respectively, rather than a single experiment with prices varying from [0%, 200%], the estimates change. There are several possible reasons for this. One is that this LLM is unable to correctly process experimental information. Another possibility is that the existing prompt still has some remaining ambiguity. For example, the LLM may make assumptions about the setting, such as the type of stores and the customers, based on the set of prices included in the experiment. Appendix F conducts additional analysis which suggests that this sensitivity to the experimental range is partly driven by the current LLMs not being able to correctly process experimental design information, and shows that newer LLMs demonstrate improved (though still imperfect) consistency across different experimental ranges. Therefore, unlike the fundamental limitations of blinding that future LLMs may not easily address, sensitivity to experimental design information could potentially be resolved through model improvements, making this a promising area for future research.
Figure 8: Unblinded design asking directly for aggregate predictions: price range comparison

### v3

@@ 3 Potential Solution: Controlling for Covariates
When working with real-world data, a common approach for causal identification is to control for covariates. If researchers have controlled for all confounding covariates that affect both the treatment and outcome, the conditional independence assumption, or ignorability, is satisfied. If sufficient variation remains after controlling for covariates, then researchers can identify the causal effect of interest. In LLM-simulated experiments, a natural starting point to control for covariates is adding detailed information about individuals and their environment in the prompt. These covariates can range from demographic information, location, preferences, budget, etc.
argyle_out_2023 introduced an approach for controlling for such covariates. They advocate for running studies on “silicon samples” of digital personas, which in their case were described by demographic variables and political beliefs that were drawn to align with nationally representative samples. This approach has been picked up by the industry and commercialized by several firms (e.g., Synthetic Users, Evidenza). Figure 3 shows that on average, controlling for covariates, such as demographic variables does indeed improve the results in our case compared to not specifying any detail. However, it does not fully resolve the issue. Web Appendix D describes the detailed prompts and covariates used, which involve creating of 500 personas for each {product,category} pair.
@@
A natural question is what constitutes a principled approach to controlling for covariates in LLM simulations and what potential pitfalls exist. When working with human observational data, researchers commonly employ the “kitchen-sink” approach, controlling for many observed pre-treatment variables because a rich set of covariates makes the conditional-independence assumption more credible by capturing confounders. This benefit is strongest with many pre-treatment variables and a large dataset that yields better empirical overlap. While LLM simulations seem well suited here—one can specify many pre-treatment covariates and simulate large samples—an important consideration is the stage at which covariates are controlled for. When working with human data, empirical researchers often control for covariates such as product characteristics and rich user history to address endogeneity during data analysis, after the data have already been collected. Hence, analyzing the data does not change the original data-generating process that has already taken place. In contrast, adding covariates via the prompt in an LLM simulation occurs at the data-generation stage. Consequently, the additional covariates can change the factors considered by a simulated individual when making decisions.
@@
For example, in the context of demand estimation, one can include contextual information in the prompt such as the price of competing products in the category, which can be used as a proxy for the type of stores that the customer is visiting and the competitive environment that the focal product is facing. However, while adding contextual variables such as competing prices in the LLM simulation may prevent the LLM from using the price treatment to infer these contextual variables, adding them may also lead the LLM-simulated customer to put a heavier emphasis on these factors when making decisions. Figure 4 shows (in the specific example of Coca-Cola) that the demand curve (elicited by simulating the choices of 500 personas) is no longer smooth after fixing such contextual variables, [^5] We fix the competing price to be the same as the regular price of the focal product. and the simulated customers’ purchase decisions follow a step function: customers will purchase if and only if the price is lower than or equal to the competing price. Web Appendix D.1 describes the detailed prompts. [^6] One possible way to fix the non-smoothness of the demand curve is to simulate a distribution of contextual variables, each with difference past and competing prices. However, this is no longer useful for a particular context, e.g., a particular store, because the simulated context deviates from the targeted context. In a real-world scenario, we should expect that even in the absence of heterogeneity in context (e.g., in a given store in a given day), there should still exists heterogeneity in consumers, such that the demand curve would be smoother. Such scenario is implausible in the real world because consumers are often inattentive and also care about other dimensions of the product characteristics. However, in LLM simulation, controlling for competing price to address endogeneity may unintentionally make the controlled covariates artificially salient to the simulated consumer and change the decision environment.
@@
Figure 4: Purchase probability, controlling for demographics and competing price: Coca-Cola example.
@@
Such phenomenon in which decision makers place more emphasis on certain aspects of a choice set or scenario than they would in real-world decisions is known in the consumer behavior literature as focalism (schkade1998does; wilson2000focalism), and constitutes a threat to ecological validity. For example, in the context of conjoint analysis, bedi2021damaged argue that when minor attributes are presented together with major ones, they tend to receive undue attention from respondents (due to focalism), leading to upward biased estimates of the importance of these minor features. While it has been documented that LLMs are prone to focalism (cheng2023compost) and focalism is not an issue that is unique to LLM experiments, the nature of the risk differs as it introduces a new set of trade-off: introducing additional covariates in the simulation may improve unconfoundedness but sacrifice ecological validity. Therefore, unlike standard experimental analysis where controlling covariates typically improve the performance, in LLM simulation, adding more covariates may degrade the performance, because it render the simulated human behavior less realistic.
To illustrate this trade-off empirically, we leverage a publicly available dataset from toubia2025twin collected through multiple survey waves containing diverse measures of consumer cognitive and psychological traits alongside demographic information. Importantly, toubia2025twin also replicated our pricing study on their panel, allowing us to relate the results of a study with the exact same design as ours to other variables collected on the same sample of consumers. We implement a sequential covariate addition procedure across twelve stages, beginning with a baseline of 14 demographic variables (age, education, income, and others), then systematically adding behavioral economics and psychological measures one at a time. This granular approach introduces individual covariates sequentially: tightwad-spendthrift score (stage 2), discount rate and present bias (stage 3), risk and loss aversion (stages 4-5), financial literacy and numeracy (stages 6-7), mental accounting (stage 8), decision-making styles (stages 9-10), environmental attitudes (stage 11), and the Big Five personality traits (stage 12), expanding from 14 to 30 total control variables.
Figure 5 tracks the mean absolute error (MAE) between actual consumer choices in toubia2025twin’s replication of our study and LLM predictions as we gradually add covariates. Appendix E provides detailed results showing the non-monotonic effects at each stage, Table E.1 reports the additional covariates added at each stage and Prompt E documents the prompt templates used. In standard experimental analyses with human data, controlling for additional covariates typically improves model performance, provided the sample size is sufficient. However, our results reveal a non-monotonic pattern: initial covariate additions improve accuracy, but beyond a threshold, further additions degrade performance. This deterioration may occurs because explicitly specifying covariates in the prompt induces focalism, in which the LLM assigns disproportionate weight to variables that consumers would typically ignore in naturalistic decision contexts, thereby compromising ecological validity despite reducing confounding.
Figure 5: Mean absolute error progression as more covariates are controlled in the simulation. The MAE is calculated by comparing the purchase probability for each (product, price) combination.
4 Core Issue: Ambiguous Prompting Strategy
So far we have empirically explored one context (demand estimation) with 40 products and one specific version of an LLM. A natural question is whether the issue we have uncovered applies generally to all LLMs and other experiment contexts. This section formalizes the key issue and shows that one key step to harness the full potential of LLM in simulating human behavior lies in avoiding ambiguous prompting strategies. Formally,
@@ 4 Core Issue: Ambiguous Prompting Strategy
Such ambiguous prompting strategies should be avoided because - for any given LLM model, no matter how advanced they are, if researchers use an ambiguous prompting strategy, then if it happens to correctly answer one question correctly, $q_{1}$, it will incorrectly answer its ambiguous counterpart, $q_{2}$. Web Appendix F.1 provides a more detailed proof.
An important step for practitioners to avoid ambiguous prompting strategy is to examine whether a specific prompt being used is open to multiple interpretations. For example, consider a simple prompt $p$ under a blinded simulation design:
Prompt 3: A simple ambiguous prompt Consider the product category: {category}. Suppose you are in a grocery store, and you see the following product in that category: {product}. The product is priced at: {price}. You decide to ___ [’purchase’ or ’not purchase’]
This prompt is ambiguous because it allows for at least two interpretations:
These two questions yield different answers. Using the do-operator (Pearl2009) to distinguish active intervention from observation, and denoting $Y$ as Purchase and $D$ as Price, the answer to $q_{1}$ is $a_{1}:P(Y|do(D))$, where $do(D)$ is randomized and independent of unobserved factors $U$ ($D\perp U$). The answer to $q_{2}$ is $a_{2}:P(Y|D)$, where $D$ can be correlated with other confounders $U$ that affect $Y$ ($D\not\perp U$), as in observational data. Since LLMs are trained primarily on observational data, it is plausible that an ambiguous question will also be interpreted in a way that is similar to observational data, hence answering the wrong question. Appendix F.2 provides a DAG illustration of the different types of data generating process that an LLM may mimic when an ambiguous prompt is used.
As discussed in Section 3, one might attempt to address this issue by adding more variables to the prompt. However, adding variables to the prompt can actually increase rather than decrease ambiguity from a causal perspective. To illustrate this multiplication of possible interpretations in practice, consider the following detailed example:
This prompt, while detailed, may be even more ambiguous. It permits more valid causal interpretations:
  1. 1.
    Price as Treatment: $P(Y|do(\text{price}),\text{health\_info},\text{sugar\_content})$
  2. 2.
    Health Information as Treatment: $P(Y|\text{price},do(\text{health\_info}),\text{sugar\_content})$: “What happens when we experimentally expose customers to health articles, holding fixed price at \$4.99?”
  3. 3.
    Sugar Content as Treatment: $P(Y|\text{price},\text{health\_info},do(\text{sugar\_content}))$: “What happens when we experimentally highlight sugar content, holding fixed price and prior health awareness?”
5 The value of unblinding
Our theoretical framework points toward a promising direction: designing unambiguous prompts by explicitly communicating the experimental design, which by definition entails unblinding. The causal inference literature (AngristPischke2009; ImbensRubin2015; athey2017econometrics) has long emphasized the importance of clearly specifying estimands and sources of variation. Following this principle, we argue that unblinded prompts should specify both the treatment being randomized and the distribution from which the treatment is drawn. We test unblinding through two approaches: first by examining out-of-the-box LLMs to establish baseline differences, and then by generalizing to fine-tuning experiments.
We begin by examining whether unblinding improves performance in standard, pre-trained language models, establishing a baseline for how prompt design affects predictions without task-specific training. To assess the robustness of the effects, we evaluate a series of GPT models spanning multiple generations, from GPT-3.5-Turbo to the latest models available at the time of writing (GPT-4.1). We operationalize the distinction between blinded and unblinded approaches through two different system prompts.
Prompt 5: Blinded System Prompt System: You, AI, are a customer. Your task is to fill in the blank ___. Return the completed information without extra text.
The unblinded condition alters the system prompt to include experimental context:
Prompt 6: Unblinded System Prompt System: You, AI, are an expert in predicting customer behavior. The customer is given a survey on their purchase decision for the {product} in {category} where the price of the product is randomly and uniformly drawn from {min_price} to {max_price}. The customer is only presented with one price and is blind to this randomization design. The customer is given the following survey. Your task is to fill in the blank ___. Return the completed information without extra text.
This design choice operationalizes our theoretical framework. The unblinded prompt explicitly communicates the data-generating process, including both the randomization mechanism and the support of the price distribution. Figure 6 reveals a consistent pattern across model generations. Unblinded prompts systematically generate more accurate predictions, with performance improvements ranging from 1% to over 60% relative to baseline. The mean absolute error (MAE) reduction persists across all models tested, including reasoning models, and this performance advantage remains substantial as models advance.
Figure 6: Gain from unblinding for various models (ordered by release date and reasoning ability). The figure shows MAE performance improvement when moving from blinded to unblinded prompts, with GPT models ordered by release date. The MAE is calculated by comparing the purchase probability for each (product, price) combination, and then the improvement is normalized by the baseline MAE of the blinded prompts. Positive values indicate superior performance for unblinded prompts.
While these out-of-the-box results demonstrate the value of unblinding, they raise a natural question: will this technique become obsolete as LLMs improve, particularly through fine-tuning? We argue that unblinding is not merely a strategy for evaluating and using pre-trained models but an important, complementary step in the model development and fine-tuning process itself.
To illustrate, we consider fine-tuning a model on actual human data. When performing fine-tuning, researchers must decide what context to provide. For instance, should they reveal the experimental origins of the data (an unblinded approach, as in Prompt 5) or choose to not disclose such context (a blinded approach, as in Prompt 5)? This choice is consequential. Even if fine-tuning helps the model learn the correlation between price and purchase, without the context of randomization, it has an insufficient basis for interpreting this correlation as causal. Explicitly stating that the price was randomized provides context; hence, it may help the model interpret the learned relationship as causal and thereby improving its potential to generalize. Appendix F.3 provides a more detailed discussion.
To empirically demonstrate this complementarity, we fine-tune models using our demand estimation survey data. We design an out-of-distribution test by performing a leave-one-category-out cross-validation. We divide the 40 products into four category groups: beverages (including juices, soft drinks, beer, and dairy drinks), refrigerated goods (including meats, dairy products, and frozen foods), snacks and bakery (including chips, cookies, coffee, candy, and ice cream), and household and pet supplies (including cleaning products, paper goods, pet food, and medications). In each fold, we fine-tune a model on three category groups and validate on the held-out fourth category group. This approach tests for true generalization, as the model must learn principles that transfer across fundamentally different product types rather than simply interpolating between similar items. This mirrors real-world applications where practitioners may be interested in forecasting scenarios that historical data has not observed before.
For each of the four cross-validation folds, we fine-tune GPT-4o-mini models under a 2 (blinded vs. unblinded prompt in fine-tuning stage) × 2 (blinded vs. unblinded prompt in evaluation stage) factorial design, varying the prompt type (blinded Prompt 5 vs. unblinded Prompt 5) used during both the fine-tuning and evaluation stages. This design allows us to disentangle the effects of unblinding at each stage. To evaluate performance, we test each model on its held-out category, generating 50 responses per product-price pair (with temperature set to 1) and calculating the purchase probability. We then measure the MAE against the true probabilities from our human survey.
Table 1 shows that fine-tuning consistently improves performance over the out-of-the-box model. However, the most accurate results are achieved when the model is both fine-tuned and evaluated with unblinded prompts. This combination reduces the MAE by an additional 15% compared to the best-performing blinded model, from 0.134 to 0.113. This finding highlights that unblinding offers a significant, nearly costless performance improvement, complementary to the more resource-intensive process of data acquisition and fine-tuning.
Table 1: MAE of Fine-Tuned Models on Survey Data
Model Evaluation Method
Blinded Unblinded
Out-of-box GPT-4o-mini 0.532 0.397
Fine-tuned on Survey (Blinded) 0.134 0.128
Fine-tuned on Survey (Unblinded) 0.130 0.113
The value of unblinding becomes even more pronounced when fine-tuning models on complex or mixed-quality datasets. While the previous exercise demonstrated unblinding’s value with a single-source high-quality dataset, practical applications often involve greater uncertainty. Developers may know a model’s general purpose—such as forecasting demand under different pricing strategies—without knowing its exact future applications. Even in a simple setting where a firm only sells two products A and B, the application can include 1) forecasting demand when A and B are priced as usual 2) adjusting the price for product A when B is priced as usual 3) adjusting the price for product B when A is priced as usual, 4) adjusting price for both product A and B, 5) predicting the introduction of a new product C when product A and B are priced as usual. To address this range of potential queries, models may be fine-tuned on diverse training data comprising both experimental datasets, each with specific treatments and covariates, and observational datasets that capture business-as-usual patterns and correlations. While observational data can answer routine questions, it may prove problematic for certain counterfactual analyses. Without additional context during fine-tuning, experimental and observational data risk becoming indistinguishable to the model. Moreover, ambiguous prompts leave the model uncertain about which dataset to leverage for specific questions. Unblinding thus enhances model robustness against irrelevant datasets that may not directly address researchers’ intended questions.
To illustrate the point, we replicate the aforementioned fine-tuning exercise, with the exception that we also added an irrelevant observational dataset that has documented how customers browse online on Amazon (berke2024open). The dataset is by construction observational and biased because it has only recorded purchases and no recorded non-purchases. Appendix G documents the exact prompt we used to incorporate the additional Amazon data in the fine-tuning process.
Table 2 shows that when this irrelevant dataset is included, the performance of models trained with blinded prompts degrades significantly. In contrast, models trained with unblinded prompts prove far more robust. By providing clear context about the data’s origin, unblinding enables the model to correctly weight the experimental data and discount the noisy observational data, preserving its predictive accuracy.
Table 2: MAE of Fine-Tuned Models on Mixed (Survey + Amazon) Data
Model Evaluation Method
Blinded Unblinded
Out-of-box GPT-4o-mini 0.532 0.397
Fine-tuned on Survey+Amazon (Blinded) 0.233 0.126
Fine-tuned on Survey+Amazon (Unblinded) 0.145 0.120